# Game-Theoretic Multiagent Reinforcement Learning

### v4

@@ 7.2 Discrete-Action Normal-Form Games
The Hedge algorithm as Eq. (7.34) achieves a regret of $\mathcal{O}(\sqrt{T\log A})$ for multi-player general-sum game, with each player having action space of size $A$[108]. EXP3 [19] extended the Hedge algorithm for a partial information game in which the player knows only the reward of the chosen action (i.e., a bandit version) and has to estimate the loss of the actions that it does not select. [69] augmented the Hedge algorithm with a tree-pruning technique based on dynamic thresholding. [209] developed Lagrangian hedging, which unifies no-regret algorithms, including both regret matching and Hedge, through a class of potential functions. Optimistic Hedge (or called OMWU) provably achieves $\mathcal{O}(N\log A\log^{4}T)$ regret for N-player general-sum games, with each player having action space of size $A$[126].
@@
$\pi^{i}_{t+1}(S,a)=\left\{\begin{array}[]{ll}\dfrac{\lfloor\mathsf{Reg}_{t}^{i}\big{(}S,a\big{)}\rfloor_{+}}{\sum_{a\in\chi(S)}\lfloor\mathsf{Reg}_{t}^{i}\big{(}S,a\big{)}\rfloor_{+}}&\text{if }\sum_{a\in\chi(S)}\lfloor\mathsf{Reg}_{t}^{i}(S,a)\rfloor_{+}>0\\ \dfrac{1}{|\chi(S)|}&\text{otherwise }\end{array}\right.$ (7.39)
@@ 7.3 Continuous-Action Normal-Form Games
Based on the recent research results above, GDA methods are notoriously flawed from three aspects. First, GDA algorithms may not converge at all [32, 379, 132], resulting in limited cycles [^2] Limited cycle is a terminology in the study of dynamical systems, which describes oscillatory systems. In game theory, an example of limit cycles in the strategy space can be found in Rock-Paper-Scissor game. in which even the time average [^3] In two-player two-action games, [488] showed that the time average payoffs would converge to a NE value if their policies do not. does not coincide with NE [369]. Second, there exist undesired stable stationary points for the GDA algorithms that are not local optima of the game [5, 369]. Third, there exist games whose equilibria are not the attractors of GDA methods at all [369]. These problems are partly caused by the intransitive dynamics (e.g., a typical intransitive game is rock-paper-scissors game) that are inherent in zero-sum games [417, 32] and the fact that each agent may have a non-smooth objective function. In fact, even in simple linear-quadratic games, the reward function cannot satisfy the smoothness condition [^4] A differentiable function is said to be smooth if the gradients of the function are continuous. globally, and the games are surprisingly not convex either [165, 369, 606].
@@
The third mainstream idea is to analyse the loss landscape from a game-theoretic perspective and design corresponding algorithms that mitigate oscillatory behaviour. Compared to the previous two mainstream ideas, which helped generate more theoretical insights than applicable algorithms, works within this stream demonstrate strong empirical improvements in training GANs. [382] investigated the game Hessian and identified that issues on the eigenvalues trigger the limited cycles. As a result, they proposed a new type of update rule based on consensus optimisation, together with a convergence guarantee to a local NE in smooth two-player zero-sum games. [5] leveraged the curvature information of the loss landscape to propose algorithms in which all stable limit points are guaranteed to be local NEs. Similarly, [370] took advantage of the differential structure of the game and constructed an algorithm for which the local NEs are the only attracting fixed points. [601] introduced a “smoothing” scheme which can be combined with GDA to stabilize the oscillation and ensure convergence to a stationary solution. [573] established new convergence results for alternating GDA and smoothed GDA under the mild assumption that the objective satisfies the Polyak-Lojasiewicz (PL) condition about one variable. [235] proposed GDA with anderson mixing (GDA-AM) which can achieve global convergence for bilinear problems under mild conditions. GAD-AM views the GDA dynamics as a fixed-point iteration and solves it using Anderson Mixing to converge to the local minimax. [412] proposed a multi-step GDA (MGDA) algorithm that finds an ${\epsilon}$–first order stationary point of the game in iterations. [315] proposed a new semi-anchoring (SA) technique for the MGDA method. This makes the MGDA method find a stationary point of a structured nonconvex-nonconcave composite minimax problem. [35] showed that a simple multi-step proximal gradient descent-ascent algorithm converges to ${\epsilon}$–first order Nash equilibrium of the min-max game. Finally, some studies have proved some characteristics of GDA. [388] studied GDA flow and proved global convergence under mild assumptions. [170] studied the role that a finite timescale separation parameter has on the learning dynamics in GDA. [541] proved that GDA dynamics can exhibit a variety of behaviors antithetical to convergence to the game theoretically meaningful min-max solution in some specific instances. While studying the properties of GDA, the results in [146] enhanced the convergence properties of its discrete-time counterpart. In addition, [129, 379] addressed the issue of limit cycling behaviour in training GANs by proposing to apply the technique of optimistic mirror descent (OMD), which is initially proposed by [441]. OMD achieves the last-iterate convergent guarantee in bilinear convex-concave games. Specifically, at each time step, OMD adjusts the gradient of that time step by considering the opponent policy at the next time step. Let $M_{t+1}$ be the predictor of the next iteration gradient [^5] In practice, it is usually set as the last iteration gradient.; we can write OMD as follows.
@@
Due to the advantages of OMD, many variants continue to improve this method, such as [379] proved that OMD converges in all coherent problems by taking an “extra-gradient” step. [545] proposed optimistic competitive gradient optimization (OCGO), an optimistic variant, for which they show convergence rate to saddle points in $\alpha$-coherent class of functions. [129] established that a variant of the widely used Gradient Descent/Ascent procedure, called Optimistic Gradient Descent/Ascent (OGDA), exhibits last-iterate convergence to saddle points in unconstrained convex-concave min-max optimization problems. [64] rediscovered a method related to OGDA, for the deterministic and the stochastic problem they showed a convergence rate of $O({1}/{k})$ and $O({1}/\sqrt{k})$, [463] proposed competitive mirror descent (CMD): a general method for solving such problems based on first order information that can be obtained by automatic differentiation. [253] proposed an accelerated stochastic mirror descent ascent (VR-SMDA) method based on the variance reduced technique. [255] proposed a distributed mirror descent algorithm for computing a Nash equilibrium and proved its convergence with suitably selected diminishing step-sizes for a strictly convex-concave cost function. [561] proposed and analyze the alternating mirror descent algorithm, in which each player takes turns to take action following the mirror descent algorithm for constrained optimization. Some work further reveals the properties of OMD, such as [13] revealed that OMD either reaches arbitrarily close to a Nash equilibrium, or it outperforms the robust price of anarchy in efficiency. [11] showed that when OMD does not reach arbitrarily close to a NE, the cumulative regret of both players is not only negative, but decays linearly with time. [190] proposed a second-order extension of the continuous-time game-theoretic mirror descent dynamics which converges to mere variationally stable states. [55] proved novel convergence results which matching the $1/{k}$ rate for the best iterate in terms of the squared operator norm recently shown for the EG for a generalized version of the OGDA. [12] found that OGD and its variants are instances of proportional integral derivative (PID) control.
@@
The other approach is extragradient (EG), EG (Korpelevich,1976) is one of the most popular methods for solving saddle point and variational in equalities problems (VIP). there remain important open questions about convergence of EG. [208] derived the first last-iterate $O({1}/{k})$ convergence rate for EG for monotone and Lipschitz VIP without any additional assumptions on the operator. [153] proposed an additional variance reduction mechanism for EG to obtain speed-ups in smooth convex games. [354] unified and established a best-convergence rate of two variants of the extragradient method for approximating a solution of a co-monotone inclusion constituted by the sum of two operators. [332] studied a class of stochastic minimax methods and develop a communication-efficient distributed stochastic extragradient algorithm, LocalAdaSEG, with an adaptive learning rate suitable for solving convex-concave minimax problems in the Parameter-Server model. [314] proposed a two-time-scale variant of the EG, named EG+, with a slow $O({1}/{k})$ rate on the squared gradient norm, where $k$ denotes the number of iterations. [106] developed provably efficient extragradient methods to find the quantal response equilibrium (QRE)—which are solutions to zero-sum two-player matrix games with entropy regularization—at a linear rate. [324] presented an analysis of the same-sample Stochastic EG (SEG) method with constant step size. [234] proposed AGE, an alternating extra-gradient method with nonlinear gradient extrapolation. It estimates the lookahead step using a nonlinear mixing of past gradient sequences. [149] presented a stochastic accelerated gradient EG (AG-EG) that achieves such a relatively mature characterization of optimality in saddle-point optimization. [164] showed that its EG algorithm reaches first an ${\eta}^{1/{\rho}}$-approximate Nash equilibrium, with ${\rho}>1$. Finally, [23] proved that EG achieves the optimal rate for a wide class of algorithms with any number of extrapolations.
@@
There are some other methods which provid us with a new perspective are not directly related to gradient method,such as [212] optimized a different objective that circumvents the min-max structure using the notion of duality gap from game theory. Convergence and stability properties of the method proposed in [460] are robust to strong interactions between the players, without adapting the step-size. [179] proposed a stochastic relaxed forward-backward (SRFB) algorithm for GANs, and we show convergence to an exact solution when an increasing number of data is available. [247] took an in-depth look at a comprehensive class of state-of-the art algorithms and prevalent heuristics in non-convex / non-concave problems. [427] investigated how adapting the reward (by adding a regularization term) of the game can give strong convergence guarantees in monotone games. [321] addressed and solved this games with a new algorithm named Stable Opponent Shaping (SOS), inheriting strong convergence guarantees from lookahead and shaping capacity from (Learning with Opponent-Learning Awareness)LOLA. To accelerate the calculation of equilibrium, [583] showed that these accelerated algorithms exhibit what we call the merging path (MP) property. [24] used matrix iteration theory to characterize acceleration in smooth games.
@@ 7.4.1 Tabular Stochastic Games
From a theoretical view, we survey the existing methods for tabular SG/MG with finite and discrete spaces, for both the state and the action. For simplicity, we denote the action space size $A=|\mathbb{A}|$ (and $B=|\mathbb{B}|$ for second player in the game), and the state space size $S=|\mathbb{S}|$, where ||$\cdot$|| ⋅ | denotes the cardinality of the countable set. The horizon of the game is $H$. The tabular case for two-player zero-sum SGs means $A$ and $B$ are finite and small.
@@ 7.5.1 Model-Free Approaches
$\pi^{i}_{T+1}(S,a)=\left\{\begin{array}[]{ll}\dfrac{\lfloor\mathsf{Reg}_{T}^{i}\big{(}S,a\big{)}\rfloor_{+}}{\sum_{a\in\chi(S)}\lfloor\mathsf{Reg}_{T}^{i}\big{(}S,a\big{)}\rfloor_{+}}&\text{if }\sum_{a\in\chi(S)}\lfloor\mathsf{Reg}_{T}^{i}(S,a)\rfloor_{+}>0\\ \dfrac{1}{|\chi(S)|}&\text{otherwise }\end{array}\right.$ (7.62)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\pi^{i,*}\in\mathbf{Br}(\pi^{-i})=\Big{\{}\arg\max_{\hat{\pi}\in\Delta(\mathbb{A}^{i})}\mathbb{E}_{\hat{\pi}^{i},\pi^{-i}}\big{[}R^{i}+\lambda\cdot\log(\hat{\pi})\big{]}\Big{\}}.$ (7.67)
@@ 7.6 Online Markov Decision Processes
$\mathsf{Reg}_{T}=\sum_{s\in\mathbb{S}}\mu^{*}(s)\sum_{t=1}^{T}\underbrace{\sum_{a\in\mathbb{A}}\Big{(}\pi^{*}(a\mid s)-\pi_{t}(a\mid s)\Big{)}Q_{t}\big{(}s,a\big{)}}_{\text{local regret in state $s$ with reward function $Q_{t}(s,\cdot)$}}.$ (7.72)
@@ 8.1 Solutions by Mathematical Programming
$\begin{array}[]{l}{\min_{\mathbf{V},\bm{\pi}}f(\mathbf{V},\bm{\pi})=\sum_{i=1}^{2}\bm{1}_{|\mathbb{S}|}^{T}\bigg{[}V^{i}-\Big{(}\mathbf{R}^{i}(\bm{\pi})+\gamma\cdot\mathbf{P}(\pi)V^{i}\Big{)}\bigg{]}}\\ \text{ s.t. }{\begin{array}[]{ll}{\text{ (a) }\pi^{2}(s)^{T}\Big{[}\mathbf{R}^{1}(s)+\gamma\cdot\sum_{s^{\prime}}\mathbf{P}(s^{\prime}|s)V^{1}(s^{\prime})\Big{]}\leq V^{1}(s)\bm{1}_{|\mathbb{A}^{1}|}^{T},\ \ \ \ \forall s\in\mathbb{S}}\\ {\text{ (b) }\Big{[}\mathbf{R}^{2}(s)+\gamma\cdot\sum_{s^{\prime}}\mathbf{P}(s^{\prime}|s)V^{2}(s^{\prime})\Big{]}\pi^{1}(s)\leq V^{2}(s)\bm{1}_{|\mathbb{A}^{2}|},\ \ \ \ \forall s\in\mathbb{S}}\\ {\text{ (c) }\pi^{1}\left(s\right)\geq\mathbf{0},\ \ \pi^{1}(s)^{T}\bm{1}_{|\mathbb{A}^{1}|}=1,\ \ \ \ \forall s\in\mathbb{S}}\\ {\text{ (d) }\pi^{2}\left(s\right)\geq\mathbf{0},\ \ \pi^{2}(s)^{T}\bm{1}_{|\mathbb{A}^{2}|}=1,\ \ \ \ \forall s\in\mathbb{S}}\end{array}}\end{array}$ (8.1)

### v5

@@ 7.2 Discrete-Action Normal-Form Games
The Hedge algorithm as Eq. (7.34) achieves a regret of $\mathcal{O}(\sqrt{T\log A})$ for multi-player general-sum game, with each player having action space of size AA[108]. EXP3 [19] extended the Hedge algorithm for a partial information game in which the player knows only the reward of the chosen action (i.e., a bandit version) and has to estimate the loss of the actions that it does not select. [69] augmented the Hedge algorithm with a tree-pruning technique based on dynamic thresholding. [209] developed Lagrangian hedging, which unifies no-regret algorithms, including both regret matching and Hedge, through a class of potential functions. Optimistic Hedge (or called OMWU) provably achieves $\mathcal{O}(N\log A\log^{4}T)$ regret for N-player general-sum games, with each player having action space of size AA[126].
@@
[…]
@@ 7.3 Continuous-Action Normal-Form Games
Based on the recent research results above, GDA methods are notoriously flawed from three aspects. First, GDA algorithms may not converge at all [32, 379, 132], resulting in limit cycles [^2] Limit cycle is a terminology in the study of dynamical systems, which describes oscillatory systems. In game theory, an example of limit cycles in the strategy space can be found in Rock-Paper-Scissor game. in which even the time average [^3] In two-player two-action games, [488] showed that the time average payoffs would converge to an NE value if their policies do not. does not coincide with NE [369]. Second, there exist undesired stable stationary points for the GDA algorithms that are not local optima of the game [5, 369]. Third, there exist games whose equilibria are not the attractors of GDA methods at all [369]. These problems are partly caused by the intransitive dynamics (e.g., a typical intransitive game is rock-paper-scissors game) that are inherent in zero-sum games [417, 32] and the fact that each agent may have a non-smooth objective function. In fact, even in simple linear-quadratic games, the reward function cannot satisfy the smoothness condition [^4] A differentiable function is said to be smooth if the gradients of the function are continuous. globally, and the games are surprisingly not convex either [165, 369, 606].
@@
The third mainstream idea is to analyse the loss landscape from a game-theoretic perspective and design corresponding algorithms that mitigate oscillatory behaviour. Compared to the previous two mainstream ideas, which helped generate more theoretical insights than applicable algorithms, works within this stream demonstrate strong empirical improvements in training GANs. [382] investigated the game Hessian and identified that issues on the eigenvalues trigger the limit cycles. As a result, they proposed a new type of update rule based on consensus optimisation, together with a convergence guarantee to a local NE in smooth two-player zero-sum games. [5] leveraged the curvature information of the loss landscape to propose algorithms in which all stable limit points are guaranteed to be local NEs. Similarly, [370] took advantage of the differential structure of the game and constructed an algorithm for which the local NEs are the only attracting fixed points. [601] introduced a “smoothing” scheme which can be combined with GDA to stabilize the oscillation and ensure convergence to a stationary solution. [573] established new convergence results for alternating GDA and smoothed GDA under the mild assumption that the objective satisfies the Polyak-Lojasiewicz (PL) condition about one variable. [235] proposed GDA with anderson mixing (GDA-AM) which can achieve global convergence for bilinear problems under mild conditions. GAD-AM views the GDA dynamics as a fixed-point iteration and solves it using Anderson Mixing to converge to the local minimax. [412] proposed a multi-step GDA (MGDA) algorithm that finds an $\epsilon$}–first order stationary point of the game in iterations. [315] proposed a new semi-anchoring (SA) technique for the MGDA method. This makes the MGDA method find a stationary point of a structured nonconvex-nonconcave composite minimax problem. [35] showed that a simple multi-step proximal gradient descent-ascent algorithm converges to $\epsilon$}–first order Nash equilibrium of the min-max game. Finally, some studies have proved some characteristics of GDA. [388] studied GDA flow and proved global convergence under mild assumptions. [170] studied the role that a finite timescale separation parameter has on the learning dynamics in GDA. [541] proved that GDA dynamics can exhibit a variety of behaviors antithetical to convergence to the game theoretically meaningful min-max solution in some specific instances. While studying the properties of GDA, the results in [146] enhanced the convergence properties of its discrete-time counterpart. In addition, [129, 379] addressed the issue of limit cycling behaviour in training GANs by proposing to apply the technique of optimistic mirror descent (OMD), which is initially proposed by [441]. OMD achieves the last-iterate convergent guarantee in bilinear convex-concave games. Specifically, at each time step, OMD adjusts the gradient of that time step by considering the opponent policy at the next time step. Let $M_{t+1}$ be the predictor of the next iteration gradient [^5] In practice, it is usually set as the last iteration gradient.; we can write OMD as follows.
@@
Due to the advantages of OMD, many variants continue to improve this method, such as [379] proved that OMD converges in all coherent problems by taking an “extra-gradient” step. [545] proposed optimistic competitive gradient optimization (OCGO), an optimistic variant, for which they show convergence rate to saddle points in $\alpha$-coherent class of functions. [129] established that a variant of the widely used Gradient Descent/Ascent procedure, called Optimistic Gradient Descent/Ascent (OGDA), exhibits last-iterate convergence to saddle points in unconstrained convex-concave min-max optimization problems. [64] rediscovered a method related to OGDA, for the deterministic and the stochastic problem they showed a convergence rate of O(1/k)O({1}/{k}) and $O({1}/\sqrt{k})$, [463] proposed competitive mirror descent (CMD): a general method for solving such problems based on first order information that can be obtained by automatic differentiation. [253] proposed an accelerated stochastic mirror descent ascent (VR-SMDA) method based on the variance reduced technique. [255] proposed a distributed mirror descent algorithm for computing a Nash equilibrium and proved its convergence with suitably selected diminishing step-sizes for a strictly convex-concave cost function. [561] proposed and analyze the alternating mirror descent algorithm, in which each player takes turns to take action following the mirror descent algorithm for constrained optimization. Some work further reveals the properties of OMD, such as [13] revealed that OMD either reaches arbitrarily close to a Nash equilibrium, or it outperforms the robust price of anarchy in efficiency. [11] showed that when OMD does not reach arbitrarily close to a NE, the cumulative regret of both players is not only negative, but decays linearly with time. [190] proposed a second-order extension of the continuous-time game-theoretic mirror descent dynamics which converges to mere variationally stable states. [55] proved novel convergence results which matching the 1/k1/{k} rate for the best iterate in terms of the squared operator norm recently shown for the EG for a generalized version of the OGDA. [12] found that OGD and its variants are instances of proportional integral derivative (PID) control.
@@
The other approach is extragradient (EG), EG (Korpelevich,1976) is one of the most popular methods for solving saddle point and variational in equalities problems (VIP). there remain important open questions about convergence of EG. [208] derived the first last-iterate O(1/k)O({1}/{k}) convergence rate for EG for monotone and Lipschitz VIP without any additional assumptions on the operator. [153] proposed an additional variance reduction mechanism for EG to obtain speed-ups in smooth convex games. [354] unified and established a best-convergence rate of two variants of the extragradient method for approximating a solution of a co-monotone inclusion constituted by the sum of two operators. [332] studied a class of stochastic minimax methods and develop a communication-efficient distributed stochastic extragradient algorithm, LocalAdaSEG, with an adaptive learning rate suitable for solving convex-concave minimax problems in the Parameter-Server model. [314] proposed a two-time-scale variant of the EG, named EG+, with a slow O(1/k)O({1}/{k}) rate on the squared gradient norm, where $k$ denotes the number of iterations. [106] developed provably efficient extragradient methods to find the quantal response equilibrium (QRE)—which are solutions to zero-sum two-player matrix games with entropy regularization—at a linear rate. [324] presented an analysis of the same-sample Stochastic EG (SEG) method with constant step size. [234] proposed AGE, an alternating extra-gradient method with nonlinear gradient extrapolation. It estimates the lookahead step using a nonlinear mixing of past gradient sequences. [149] presented a stochastic accelerated gradient EG (AG-EG) that achieves such a relatively mature characterization of optimality in saddle-point optimization. [164] showed that its EG algorithm reaches first an ${\eta}^{1/{\rho}}$-approximate Nash equilibrium, with ${\rho}>1$. Finally, [23] proved that EG achieves the optimal rate for a wide class of algorithms with any number of extrapolations.
@@
There are some other methods that provide us with a new perspective. For example, [212] optimized a different objective that circumvents the min-max structure using the notion of duality gap from game theory. The convergence and stability properties of the method proposed in [460] remain robust under strong interactions between players, even without adapting the step size. [179] proposed a stochastic relaxed forward-backward (SRFB) algorithm for GANs, and demonstrated convergence to an exact solution as more data becomes available. [247] conducted an in-depth study of a comprehensive class of state-of-the-art algorithms and prevalent heuristics for non-convex / non-concave problems. [427] investigated how adapting the reward—by adding a regularization term—can yield strong convergence guarantees in monotone games. [321] addressed and solved these games using a new algorithm called Stable Opponent Shaping (SOS), which inherits strong convergence guarantees from lookahead and the shaping capacity from Learning with Opponent-Learning Awareness (LOLA). To accelerate the calculation of equilibrium, [583] showed that these accelerated algorithms exhibit what we call the merging path (MP) property. [24] used matrix iteration theory to characterize acceleration in smooth games.
@@ 7.4.1 Tabular Stochastic Games
From a theoretical view, we survey the existing methods for tabular SG/MG with finite and discrete spaces, for both the state and the action. For simplicity, we denote the action space size $A=|\mathbb{A}|$ (and $B=|\mathbb{B}|$ for second player in the game), and the state space size $S=|\mathbb{S}|$, where $|\cdot|$ denotes the cardinality of the countable set. The horizon of the game is $H$. The tabular case for two-player zero-sum SGs means $A$ and $B$ are finite and small.
@@ 7.5.1 Model-Free Approaches
[…]
@@
1:Initialise: the “high-level” policy set $\mathbb{S}=\prod_{i\in\mathcal{N}}\mathbb{S}^{i}$, the meta-game payoff ${\bm{\mathsfit{M}}},\forall S\in\mathbb{S}$, and meta-policy ${\bm{\pi}}^{i}=\operatorname{UNIFORM}(\mathbb{S}^{i})$.
@@
4:   Compute the meta-policy ${\bm{\pi}}_{t}$ by meta-game solver $\mathcal{S}({\bm{\mathsfit{M}}}_{t})$.
@@
6:   Expand $\mathbb{S}_{t+1}^{i}\leftarrow\mathbb{S}_{t}^{i}\cup\{S_{t}^{i}\}$ and update meta-payoff ${\bm{\mathsfit{M}}}_{t+1}$.
@@
The double oracle algorithm (DO)[377] iteratively builds a meta-game by computing a NE of the meta-game and then adding best responses for both players to the opponent’s meta-distribution. Policy Space Response Oracles (PSRO)[408] is a direct extension of double oracle [377] that uses an RL subroutine as an approximate best response. Specifically, one can write PSRO and its variations in Algorithm 1, which essentially involves an iterative two-step process of solving for the meta-policy first (e.g., Nash over the meta-game), and then based on the meta-policy, finding a new better-performing policy, against the opponent’s current meta-policy, to augment the existing population. The meta-policy solver, denoted as $\mathcal{S}(\cdot)$, computes a joint meta-policy profile $\bm{\pi}$} based on the current payoff $\bm{\mathsfit{M}}$} where different solution concepts can be adopted (e.g., NE). Finding a new policy is equivalent to solving a single-player optimisation problem given opponents’ policy sets $\mathbb{S}^{-i}$ and meta-policies ${\bm{\pi}}^{-i}$, which are fixed and known. One can regard a new policy as given by an Oracle, denoted by $\mathcal{O}$. In two-player zero-sum cases, an oracle represents $\mathcal{O}^{1}({\bm{\pi}}^{2})=\{S^{1}:\sum_{S^{2}\in\mathbb{S}^{2}}{\bm{\pi}}^{2}(S^{2})\cdot\phi(S^{1},S^{2})>0\}$. Finally, after a new policy is found, the payoff table $\bm{\mathsfit{M}}$} is expanded, and the missing entries are filled by running new game simulations. The above two-step process loops over each player at each iteration, and it terminates if no new policies can be found for any players.
@@
[…]
@@ 7.6 Online Markov Decision Processes
[…]
@@ 8.1 Solutions by Mathematical Programming
[…]